Two paper hangers work at different rates. The first worker can wallpaper a wall in 2 hours. The second takes 3 hours to wallpaper the same wall. How long would it take the two workers working together to paper the wall?
Answer by CharlesG2(834)

Two paper hangers work at different rates. The first worker can wallpaper a wall in 2 hours. The second takes 3 hours to wallpaper the same wall. How long would it take the two workers working together to paper the wall?

total time = (1 job)/((1 job)/(time for A) + (1 job)/(time for B))
units: hrs = job/(job/hrs + job/hrs) --> hrs = job * hrs/job = hrs
time = 1/(1/2 + 1/3)
time = 1/(3/6 + 2/6)
time = 1/(5/6)
time = 6/5 hrs
time = 1.2 hrs working together to paper the wall